In Maryland, only three goods are produced domestically - popcorns, movie tickets, and diet drinks. The following table shows the price and quantities of these goods produced in the year 2018, 2019, 2020. 2018 is the base year



The market bundle for a typical family is 3 popcorns, 3 movie tickets and 5 diet drinks.



Calculate Nominal and Real GDP in all three years.


Calculate real economic growth rate in 2020


Calculate CPI in all three years.


Calculate inflation rate in 2020




2019



2020



2021





P



Q



P



Q



P



Q



Popcorn



1



500



1



600



1.05



590



Movie ticket



5



300



10



200



10.5



210



Diet drink



0.7



300



0.8



400



0.75



420

1
Expert's answer
2021-11-07T19:52:27-0500



a)Nominal and Real GDP in all three years.


Nominal GDP=(Units Produced in a Year)"\times" (Price in a Year)

For 2018: ="(500\times 1.00)+(300\times5.00)+(300\times 0.70)=2210"

For 2019: ="(600\times 1.00)+(200\times10.00)+(400\times 0.80)=2920"

For 2020: ="(590\times 1.05)+(210\times5.00)+(420\times 0.75)=3045"


Real GDP:

is equal to the sum of the base year price×current year quantity of all the goods.

For 2018: ="(500\times 1.00)+(300\times5.00)+(300\times 0.70)=2210"

For 2019: ="(600\times 1.00)+(200\times5.00)+(400\times 0.70)=1880"

For 2020: ="(590\times 1.00)+(210\times5.00)+(420\times 0.70)=1934"


b)Real economic growth rate in 2020

=((Real GDP"_{20}" -Real GDP"_{19}" )"\div" Real GDP"_{19}" ))"\times" 100


="\frac{1934-1880}{1880}\times 100=2.87\%"


c)Calculate CPI in all three years.

CPI= (Current year cost "\div" Base year cost) "\times" 100


CPI"_{18}=\frac{(5\times1.00)+(3\times5.00)+(3\times0.7)}{(5\times1.00)+(3\times5.00)+(3\times0.7)}\times100=100"


CPI"_{19}=\frac{(5\times1.00)+(3\times10.00)+(3\times0.8)}{(5\times1.00)+(3\times5.00)+(3\times0.7)}\times100=169.2"


CPI"_{20}=\frac{(5\times1.05)+(3\times10.05)+(3\times0.75)}{(5\times1.00)+(3\times5.00)+(3\times0.7)}\times100=176.5"


d)Inflation rate in 2020


"\\frac{CPI_{20}\u200b -CPI_{19}}{CPI_{19}}\\times100=\\frac{176.5-169.2}{169.2}\\times100=4.3\\%"